# Multi-band, Multi-epoch Photometry of the Spot-crossing System TOI-3884: Refined System Geometry and Spot Properties

### v1

@@ Abstract
Spot-crossing transits offer a unique opportunity to probe spot properties such as temperature, size, and surface distribution. TOI-3884 is a rare system in which spot-crossing features are persistently observed during every transit. This is due to its unusual configuration: a polar-orbiting super-Neptune transits a pole-on mid-M dwarf, repeatedly crossing a polar spot. However, previous studies have reported discrepant values in key system parameters, such as stellar inclination and obliquity. To address this, we conducted multi-band, multi-epoch transit observations of TOI-3884b using the MuSCAT instrument series, along with photometric monitoring with the LCO 1m telescopes/Sinistro. We detected time-dependent variations in the spot-crossing signals, indicating that the spot is not exactly on the pole. From the monitoring data, we measured a stellar rotation period of $11.043_{-0.053}^{+0.054}$ days with a modulation amplitude of $\sim$5in the $r$-band, consistent with the time variability in the spot-crossing features. Our analysis reconciles previous discrepancies and improves the constraints on the parameters of the system geometry ($i_{\star}=40.1^{+2.0}_{-1.2}$ deg and $\lambda=139.0^{+9.0}_{-3.7}$ deg) and those of the spot properties (spot radius of $0.425_{-0.011}^{+0.018}\,R_{\star}$ and spot-photosphere temperature difference of $200_{-9}^{+11}$ K). These results provide a critical context for interpreting upcoming transmission spectroscopy of TOI-3884 b, as well as yielding new insights into the magnetic activity and spin-orbit geometry of M dwarfs.
@@ 1 Introduction
The TOI-3884 (LSPM J1206+1230, Table 1) system offers a unique opportunity to characterize the stellar surface via spot-crossing observations. The system consists of a bright M4-dwarf host with a transiting super-Neptune, TOI-3884 b. This system is particularly unusual because the planet’s transits persistently exhibit spot-crossing signals. This is best explained by a polar-orbiting planet that transits in front of a large spot located near the stellar pole, with the system viewed nearly pole-on. The system was found and confirmed in J. M. Almenara et al. (2022, A22 hereafter), and independently characterized by J. E. Libby-Roberts et al. (2023, L23 hereafter).
@@ 4 Transit light curve Analyses
From the MCMC results, we found that the observed spot-crossing light curves can be explained by several distinct solutions. These solutions broadly fall into two groups: one with a stellar rotation period of $\sim 11.04$ days and another of $\sim 11.5$ days. The solutions near 11.5 days exhibited log-probability values that were lower by more than 50 compared to those near 11.04 days, indicating significantly worse fits. However, some chains became trapped in the less probable 11.5-day region. Therefore, we also imposed a uniform prior of $\mathcal{U}(10.8,11.3)$ on $P_{\rm rot}$, restricting it to within $5\sigma$ of the value derived from the rotational modulation. After applying this prior and rerunning the MCMC, the chains successfully converged.
@@
Spot latitude $\theta_{\rm spot}$ [deg] $\mathcal{U}$(-90, 90) $76.8_{-2.5}^{+4.6}$
Spot longitude $\varphi_{\rm spot}$ [deg] $\mathcal{U}$(0, 360) $98_{-6}^{+14}$
@@
Stellar inclination $i_{\star}$ [deg] $\mathcal{U}$(0, 90) $40.1_{-1.2}^{+2.0}$
Projected obliquity $\lambda$ [deg] $\mathcal{U}$(-180, 180) $139.0_{-3.7}^{+9.0}$
@@
Figure 3: Observed transit light curves of TOI-3884b (top to bottom: three transits observed with MuSCAT3 or MuSCAT4; left to right: $g$-, $r$-, $i$-, and $z$-band) and the corresponding best-fit models (black lines). The error bars in the datapoints are scaled to make reduced $\chi^{2}$ values equal to 1. To the right of each light curve, the spot distribution on the stellar surface at the corresponding epoch is shown, as modeled with fleck. The x mark in the figure shows the location of the stellar pole, the dotted lines show the stellar equator, and the regions inside dashed lines indicate the transit chord.
@@ 5 Discussion
Spot latitudeb $\theta_{\rm spot}$ [deg] 90 (fixed) - $76.8_{-2.5}^{+4.6}$
Spot longitudeb $\varphi_{\rm spot}$ [deg] - - $98_{-6}^{+14}$
@@
Stellar inclination $i_{\star}$ [deg] $47.0\pm 8.5$ $25\pm 5$ $40.1_{-1.2}^{+2.0}$
Projected obliquity $\lambda$ [deg] $151\pm 11$ $75\pm 10$ $139.0_{-3.7}^{+9.0}$
True obliquity $\Psi$ [deg] $130\pm 12$ $84\pm 5$c $118.1_{-2.3}^{+5.6}$
@@ 5.6 Geometry of the TOI-3884 System
Our analysis confirms the previously suggested pole-on orbital configuration of TOI-3884 b and updates the relevant system parameters. The derived stellar inclination is $i_{\star}=40.1_{-1.2}^{+2.0}$ deg and the projected obliquity is $\lambda=139.0_{-3.7}^{+9.0}$ deg, closer to the results of A22 than those of L23 (Table 4). If a stellar rotation period of 11.5 days is adopted instead, the solution becomes more consistent with that of L23. Distinguishing between these degenerate solutions would have been difficult without an accurate measurement of the stellar rotation period.
Using the derived values of $i_{\star}$, $i_{\rm orb}$, and $\lambda$, we calculate the true obliquity $\Psi=118.1_{-2.3}^{+5.6}$ deg following the geometric relation (see D. C. Fabrycky & J. N. Winn, 2009):
@@
This places TOI-3884 b among the small population of Neptune-sized planets known to have polar orbits (e.g., E. M. Louden & S. C. Millholland, 2024).
Figure 8 presents the distribution of obliquities for systems where $\lambda$ or $\Psi$ has been measured. Planets orbiting M dwarfs with similar obliquities include GJ 436b (V. Bourrier et al., 2022) and GJ 3470b (G. Stefànsson et al., 2022), both of which were confirmed to have polar orbits through Rossiter–McLaughlin (RM) effect measurements. However, their host stars are earlier-type and more slowly rotating than TOI-3884, and it remains unclear whether the same mechanisms are responsible for their polar orbits.
@@ 6 Conclusion
  • From our transit fitting, we derive a stellar inclination of $i_{\star}\sim 40$ deg, a projected obliquity of $\lambda\sim 139$ deg, and a true obliquity of $\Psi\sim 118$ deg, confirming the planet’s polar orbit configuration.
@@
MM is supported by JSPS Grant-in-Aid for JSPS Fellows Grant Number JP24KJ0241, and JSPS KAKENHI Grant Number JP24K17083. This work is also supported by JSPS KAKENHI Grant Number JP24H00017, JP24K00689, JP25K17450, JP24K17082, JP24H00248, and JSPS Bilateral Program Number JPJSBP120249910. This work is partly supported by JST SPRING, Grant Number JPMJSP2108. This work makes use of observations made with the MuSCAT3/4 instruments, developed by the Astrobiology Center (ABC) in Japan, the University of Tokyo, and Las Cumbres Observatory (LCOGT). MuSCAT3 was developed with financial support by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), and is located at the Faulkes Telescope North on Maui, HI (USA), operated by LCOGT. MuSCAT4 was developed with financial support provided by the Heising-Simons Foundation (grant 2022-3611), JST grant number JPMJCR1761, and the ABC in Japan, and is located at the Faulkes Telescope South at Siding Spring Observatory (Australia), operated by LCOGT. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Funding for the TESS mission is provided by NASA’s Science Mission Directorate. All the TESS data used in this paper can be found in MAST: http://dx.doi.org/10.17909/t9-nmc8-f686 (catalog 10.17909/t9-nmc8-f686). This work makes use of observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The authors used ChatGPT (version GPT-4), a language model developed by OpenAI, to help improve the grammar and clarity of the manuscript.
@@ Appendix A Comparison of posterior distributions from transit fitting
To investigate these two peaks, we divided the converged MCMC samples into two groups based on the spot latitude: group 1 includes samples with a spot latitude greater than 78 deg, and group 2 includes those with lower values. Figure 9 presents the corner plots of the posterior distributions for each group. From each group, we extracted the median value of each parameter as the best-fit solution.
@@
Figure 9: Corner plot showing the posterior distributions of the fitted parameters from the transit analysis. The plot includes a subset of the posteriors, omitting the 12 additional parameters ($k$, $q_{1}$, $q_{2}$ for each band) that were also sampled. Blue and orange contours represent posterior samples from group 1 (spot latitude $\theta_{\rm spot}>78$ deg) and group 2 ($\leq 78$ deg), respectively. The apparent overlap in the $\theta_{\rm spot}$ contours is due to smoothing in the visualization and does not indicate actual mixing. The blue and orange lines indicate the best-fit parameter values for each group, respectively. The green solid line in each histogram marks the best-fit value from the full posterior.
@@ Appendix C Reanalysis of TESS light curves
The priors and posteriors for the TESS transit fit are summarized in Table 5, and the best-fit model light curves are shown in Figure 14. These results suggest a spot latitude of $\sim$88 deg, consistent with a nearly exact polar spot and the absence of rotational modulation in the TESS light curves.
@@
Spot latitude $\theta_{\rm spot}$ [deg] $\mathcal{U}$(-90, 90) $88.1_{-1.9}^{+1.3}$
Spot longitude $\varphi_{\rm spot}$ [deg] $\mathcal{U}$(0, 360) $137_{-74}^{+91}$
@@
Stellar inclination $i_{\star}$ [deg] $\mathcal{U}$(0, 90)$\times$ $\mathcal{N}$(40.1, 2.0) $38.9\pm 1.9$
Projected obliquity $\lambda$ [deg] $\mathcal{U}$(-180, 180)$\times$ $\mathcal{N}$(139.0, 9.0) $153.8_{-5.6}^{+4.2}$

### v2

@@ Abstract
Spot-crossing transits offer a unique opportunity to probe spot properties such as temperature, size, and surface distribution. TOI-3884 is a rare system in which spot-crossing features are persistently observed during every transit. This is due to its unusual configuration: a nearly poler orbit super-Neptune transits a pole-on mid-M dwarf, repeatedly crossing a polar spot. However, previous studies have reported discrepant values in key system parameters, such as stellar inclination and obliquity. To address this, we conducted multi-band, multi-epoch transit observations of TOI-3884b using the MuSCAT instrument series, along with photometric monitoring with the LCO 1m telescopes/Sinistro. We detected time-dependent variations in the spot-crossing signals, indicating that the spot is not exactly on the pole. From the monitoring data, we measured a stellar rotation period of $11.043_{-0.053}^{+0.054}$ days with a modulation amplitude of $\sim$5in the $r$-band, consistent with the time variability in the spot-crossing features. Our analysis reconciles previous discrepancies and improves the constraints on the parameters of the system geometry ($i_{\star}=139.9_{-2.0}^{+1.2}$ deg and $\lambda=41.0_{-9.0}^{+3.7}$ deg) and those of the spot properties (spot radius of $0.425_{-0.011}^{+0.018}\,R_{\star}$ and spot-photosphere temperature difference of $200_{-9}^{+11}$ K). These results provide a critical context for interpreting upcoming transmission spectroscopy of TOI-3884 b, as well as yielding new insights into the magnetic activity and spin-orbit geometry of M dwarfs.
@@ 1 Introduction
The TOI-3884 (LSPM J1206+1230, Table 1) system offers a unique opportunity to characterize the stellar surface via spot-crossing observations. The system consists of a bright M4-dwarf host with a transiting super-Neptune, TOI-3884 b. This system is particularly unusual because the planet’s transits persistently exhibit spot-crossing signals. This is best explained by a highly misaligned planet that transits in front of a large spot located near the stellar pole, with the system viewed nearly pole-on. The system was found and confirmed in J. M. Almenara et al. (2022, A22 hereafter), and independently characterized by J. E. Libby-Roberts et al. (2023, L23 hereafter).
@@ 4 Transit light curve Analyses
From the MCMC results, we found that the observed spot-crossing light curves can be explained by several distinct solutions. In particular, the stellar inclination $i_{\star}$ clusters around two regions: approximately $-40$ deg and $140$ deg. These correspond to whether the stellar rotation is counterclockwise or clockwise as seen from the observer’s perspective—that is, whether the visible pole is the north or south pole.
The solution with $i_{\star}\sim-40$ deg is associated with a stellar rotation period of approximately 11.25 days. In contrast, the solution with $i_{\star}\sim 140$ deg further splits into two groups: one with $P_{\rm rot}\sim 11.04$ days and the other with $P_{\rm rot}\sim 11.5$ days. The solutions near 11.25 and 11.5 days exhibited log-probability values more than 30 and 50 lower, respectively, than those near 11.04 days, indicating significantly worse fits.
To eliminate these less probable configurations, we imposed a uniform prior of $\mathcal{U}(90,180)$ on $i_{\star}$, and $\mathcal{U}(10.8,11.3)$ on $P_{\rm rot}$, the latter restricting the range to within $5\sigma$ of the value inferred from the rotational modulation. After applying these priors and rerunning the MCMC, the chains successfully converged.
@@
Spot latitude $\theta_{\rm spot}$ [deg] $\mathcal{U}$(-90, 90) $-76.8_{-4.6}^{+2.5}$
Spot longitude $\varphi_{\rm spot}$ [deg] $\mathcal{U}$(0, 360) $262_{-14}^{+6}$
@@
Stellar inclination $i_{\star}$ [deg] $\mathcal{U}$(90, 180) $139.9_{-2.0}^{+1.2}$
Projected obliquity $\lambda$ [deg] $\mathcal{U}$(-180, 180) $41.0_{-9.0}^{+3.7}$
@@
Figure 3: Observed transit light curves of TOI-3884b (top to bottom: three transits observed with MuSCAT3 or MuSCAT4; left to right: $g$-, $r$-, $i$-, and $z$-band) and the corresponding best-fit models (black lines). The error bars in the datapoints are scaled to make reduced $\chi^{2}$ values equal to 1. To the right of each light curve, the spot distribution on the stellar surface at the corresponding epoch is shown, as modeled with fleck. The x mark in the figure shows the location of the stellar south pole, the dotted lines show the stellar equator, and the regions inside dashed lines indicate the transit chord.
@@ 5 Discussion
Spot latitudeb $\theta_{\rm spot}$ [deg] 90 (fixed) - $-76.8_{-4.6}^{+2.5}$
Spot longitudeb $\varphi_{\rm spot}$ [deg] - - $262_{-14}^{+6}$
@@
Stellar inclination $i_{\star}$ [deg] $47.0\pm 8.5$ $25\pm 5$ $139.9_{-2.0}^{+1.2}$
Projected obliquity $\lambda$ [deg] $151\pm 11$ $75\pm 10$ $41.0_{-9.0}^{+3.7}$
True obliquity $\Psi$ [deg] $50\pm 12$ or $130\pm 12$ $84\pm 5$c $61.9_{-5.6}^{+2.3}$
@@ 5.6 Geometry of the TOI-3884 System
Our analysis confirms the previously suggested nearly polar-orbit configuration of TOI-3884 b and updates the relevant system parameters. The derived stellar inclination is $i_{\star}=139.9_{-2.0}^{+1.2}$ deg and the projected obliquity is $\lambda=41.0_{-9.0}^{+3.7}$ deg. This result, when ignoring the direction of stellar rotation, corresponds to a solution with an absolute stellar inclination of $\sim$40 deg, and is therefore closer to the results of A22, than those of L23 (Table 4). If a stellar rotation period of 11.5 days is adopted instead, the solution becomes more consistent with that of L23. Distinguishing between these degenerate solutions would have been difficult without an accurate measurement of the stellar rotation period.
Using the derived values of $i_{\star}$, $i_{\rm orb}$, and $\lambda$, we calculate the true obliquity $\Psi=61.9_{-5.6}^{+2.3}$ deg following the geometric relation (see D. C. Fabrycky & J. N. Winn, 2009):
@@
This places TOI-3884 b among the small population of Neptune-sized planets known to have close to polar orbits (e.g., E. M. Louden & S. C. Millholland, 2024).
Figure 8 presents the distribution of obliquities for systems where $\lambda$ or $\Psi$ has been measured. Planets orbiting M dwarfs with similar obliquities include GJ 436b (V. Bourrier et al., 2022) and GJ 3470b (G. Stefànsson et al., 2022), both of which were confirmed to have polar orbits through Rossiter–McLaughlin (RM) effect measurements. However, their host stars are earlier-type and more slowly rotating than TOI-3884, and it remains unclear whether the same mechanisms are responsible for their orbital misalignment.
@@ 6 Conclusion
  • From our transit fitting, we derive a stellar inclination of $i_{\star}\sim 140$ deg, a projected obliquity of $\lambda\sim 41$ deg, and a true obliquity of $\Psi\sim 62$ deg, confirming the planet’s highly misaligned configuration.
@@
After posting this manuscript on arXiv, we realized that the direction of the stellar inclination vector had been incorrectly defined. This affected several inferred parameters and descriptions, which have since been updated accordingly. We are deeply grateful to Dr. Patrick Tamburo for pointing out this important issue. We also thank the anonymous referee for their insightful suggestions, and the editor for efficiently handling the manuscript. MM is supported by JSPS Grant-in-Aid for JSPS Fellows Grant Number JP24KJ0241, and JSPS KAKENHI Grant Number JP24K17083. This work is also supported by JSPS KAKENHI Grant Number JP24H00017, JP24K00689, JP25K17450, JP24K17082, JP24H00248, and JSPS Bilateral Program Number JPJSBP120249910. This work is partly supported by JST SPRING, Grant Number JPMJSP2108. This work makes use of observations made with the MuSCAT3/4 instruments, developed by the Astrobiology Center (ABC) in Japan, the University of Tokyo, and Las Cumbres Observatory (LCOGT). MuSCAT3 was developed with financial support by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), and is located at the Faulkes Telescope North on Maui, HI (USA), operated by LCOGT. MuSCAT4 was developed with financial support provided by the Heising-Simons Foundation (grant 2022-3611), JST grant number JPMJCR1761, and the ABC in Japan, and is located at the Faulkes Telescope South at Siding Spring Observatory (Australia), operated by LCOGT. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Funding for the TESS mission is provided by NASA’s Science Mission Directorate. All the TESS data used in this paper can be found in MAST: http://dx.doi.org/10.17909/t9-nmc8-f686 (catalog 10.17909/t9-nmc8-f686). This work makes use of observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The authors used ChatGPT (version GPT-4), a language model developed by OpenAI, to help improve the grammar and clarity of the manuscript.
@@ Appendix A Comparison of posterior distributions from transit fitting
To investigate these two peaks, we divided the converged MCMC samples into two groups based on the spot latitude: group 1 includes samples with a spot latitude smaller than -78 deg, and group 2 includes those with higher values. Figure 9 presents the corner plots of the posterior distributions for each group. From each group, we extracted the median value of each parameter as the best-fit solution.
@@
Figure 9: Corner plot showing the posterior distributions of the fitted parameters from the transit analysis. The plot includes a subset of the posteriors, omitting the 12 additional parameters ($k$, $q_{1}$, $q_{2}$ for each band) that were also sampled. Blue and orange contours represent posterior samples from group 1 (spot latitude $\theta_{\rm spot}\leq-78$ deg) and group 2 ($>-78$ deg), respectively. The apparent overlap in the $\theta_{\rm spot}$ contours is due to smoothing in the visualization and does not indicate actual mixing. The blue and orange lines indicate the best-fit parameter values for each group, respectively. The green solid line in each histogram marks the best-fit value from the full posterior.
@@ Appendix C Reanalysis of TESS light curves
The priors and posteriors for the TESS transit fit are summarized in Table 5, and the best-fit model light curves are shown in Figure 14. These results suggest a spot latitude of $\sim$ -88 deg, consistent with a nearly exact polar spot and the absence of rotational modulation in the TESS light curves.
@@
Spot latitude $\theta_{\rm spot}$ [deg] $\mathcal{U}$(-90, 90) $-88.1_{-1.3}^{+1.9}$
Spot longitude $\varphi_{\rm spot}$ [deg] $\mathcal{U}$(0, 360) $223_{-91}^{+74}$
@@
Stellar inclination $i_{\star}$ [deg] $\mathcal{U}$(0, 90)$\times$ $\mathcal{N}$(139.9, 2.0) $141.1\pm 1.9$
Projected obliquity $\lambda$ [deg] $\mathcal{U}$(-180, 180)$\times$ $\mathcal{N}$(41.0, 9.0) $26.2_{-4.2}^{+5.6}$